OpenStudy (anonymous):

A miscope slide show 45 red blood cells out of 75 blood cells . How many red blood cells would be expected in a sample of the same blood that has 450???

OpenStudy (lasttccasey):

Well 45/75 is to x/450 Solve for x

OpenStudy (anonymous):

Take the ratio 45 : 75 Then you need to find what number multiplies by 75 to give 450 Multiply the 45 by the number and you'll have scaled your ratio to some x : 450

OpenStudy (lasttccasey):

45/75=x/450

OpenStudy (arabpride):

45/75 = 0.6......Then multiply 450*0.6 = 270 red blood cells